Comment epi2 SNF 2007-05-16 12:18:15: Dome update

eenriquez at eenriquez at
Wed May 16 18:00:36 PDT 2007

Mike D. cleaned dome in HF/Nitric.  Reassembled chamber with no problems.  Last chamber leak rate test was 6.6 mT/min; need to be less than 5 mT/min.  Will leave in Hot idle over night.  Maurice will run a qual in the morning.

More information about the epi2-pcs mailing list